# Learning to Imitate Spatial Organization in Multi-robot Systems

### v1

@@ Abstract
Understanding collective behavior and how it evolves is important to ensure that robot swarms can be trusted in a shared environment. One way to understand the behavior of the swarm is through collective behavior reconstruction using prior demonstrations. Existing approaches often require access to the swarm controller which may not be available. We reconstruct collective behaviors in distinct swarm scenarios involving shared environments without using swarm controller information. We achieve this by transforming prior demonstrations into features that sufficiently describe multi-agent interactions before behavior reconstruction with multi-agent generative adversarial imitation learning (MA-GAIL). We show that our approach outperforms existing algorithms in all investigated swarm scenarios, and can be used to observe and reconstruct a swarm’s behavior for further analysis and testing, which might be impractical or undesirable on the original robot swarm.
@@ I INTRODUCTION
Swarm robotics and its applications are transitioning from the laboratory to the real world [1, 2], and it is expected to lead to the large-scale deployment of multiple robots in environments occupied by humans and other types of robots. For the robot-robot and human-robot interactions to be seamless, robot swarms must be safe and trustworthy [3].
@@
Ensuring the safe and trustworthy behavior of the swarms in shared environments requires a precise and continuous knowledge of their collective behavior and how it evolves. In scenarios where the swarm controllers are directly accessible, collective behavior can be easily modeled using the controllers [4]. However, in practical swarm scenarios, swarm controllers may not be accessible for various reasons (e.g., extracting controllers from natural swarms). In robot swarms, access to controllers may be restricted or impractical because some robots are remotely operated or the controller information is intentionally encrypted or hidden for strategic or privacy reasons [5, 6]. To this end, understanding the collective behavior of a swarm and modeling its dynamics without using swarm controller information requires thorough research.
@@
Collective behavior reconstruction and recognition are two established methods for modeling swarm dynamics or explaining collective behavior [7]. Collective behavior reconstruction can be model-based or data-driven. In model-based approaches, behavior reconstruction is achieved using a mathematical or regression model [8, 9]. In data-driven approaches, multi-agent interactions are statistically extracted or learned from prior demonstrations to reproduce observed behavior. Recently, data-driven approaches leverage imitation learning (IL) algorithms such as inverse reinforcement learning (IRL) and generative adversarial imitation learning (GAIL) for improved reconstruction accuracy in swarm scenarios modeled as multi-agent systems [10, 11, 12, 13, 14, 15]. Genetic programming and graphical neural networks have also been used for data-driven behavior reconstruction with swarm controllers extracted from video demonstrations [16] and swarm behavior prediction [17]. For recognition of collective behaviour, binary classification of observed behavior as defined or undefined collective behavior is a common approach [18, 19, 20]. This approach does not scale, as each swarm scenario requires a unique classifier. Multinomial classification addresses this and has been achieved in closely related swarming scenarios using predefined and learned multi-agent interaction quantifiers [21, 22, 23].
@@
Existing data-driven behavior reconstruction approaches extract multi-agent interactions from expert demonstrations without capturing swarm-environment interactions. As a result, the recovered multi-agent interactions cannot accurately reconstruct or predict expert behavior. While [17] address this in their work, their approach relied on learned extraction of multi-agent interactions and it is difficult to explain how the robots interact with the environment.
@@
In this work, we investigate the reconstruction of collective behavior in practical swarm scenarios involving shared environments without using swarm controller information. We consider three common swarm robotic scenarios: aggregation, homing, and obstacle avoidance. We model these scenarios as single-objective swarm scenarios where swarming agents interact with each other and the environment while completing the swarm objective. We generate informed and explainable multi-agent interactions through feature transformation of expert demonstrations and, achieve near-optimal behavior reconstruction using multi-agent GAIL. We show that even when learner robots are initialized from unforeseen states, they perform similarly to the expert robots in all investigated scenarios.
@@ II RELATED WORKS
IRL has been extensively used in literature for collective behavior reconstruction because it can recover the underlying reward functions while reproducing expert behavior. Šošić et al. [11] reconstructed the behavior of a homogeneous swarm by assuming that all agents are interchangeable and share a central reward function, thereby reducing the problem to a single-agent IRL. Another study [12] reconstructed the collective behavior observed in a flock of pigeons by recovering individual reward functions for each pigeon. This individualistic approach exposed the multi-agent interactions in the flock and allowed the researchers to model the leader-follower hierarchy. In [24], a similar individualistic reward function approach was used to evolve the robot controller using IRL by manually specifying the desired goal location or the path. Other studies [13, 10] extended the IRL framework to multi-agent IRL to simultaneously recover the reward functions of several agents. Liu et al. [14], however, integrated GAIL with population-based training for collective behavior reconstruction. Aside IRL and GAIL, other machine learning methods have been used to reconstruct collective behavior. In [16], genetic programming was used to extract explainable controllers from video demonstrations of collective behavior with $8$ predefined swarm interaction metrics defining the fitness measure. Zhou et al. [17] used graphical neural networks to imitate the behavior of expert robots and predict trajectories. Their approach considers swarm-environment interactions and modeled robots and environmental entities as graph nodes, but does not provide explainable interactions or controllers. Most of these works demonstrate behavior reconstruction in a single scenario or multiple similar scenarios (e.g. swarming and schooling). However, Yu et al. [15] leveraged Adversarial Imitation Learning with parameter sharing (PS-AIRL) for behavior reconstruction in distinct swarming scenarios. Their approach focused on homogeneous biological swarms and did not consider swarm-environment interactions. Furthermore, they require access to the original swarm controllers which is rarely available in practical swarm scenarios. In contrast, our approach considers swarm-environment interactions and reconstructs expert behavior without accessing robot controllers. We also generate informed multi-agent interactions that can be used to explain swarm behavior.
@@ IV-B Policy Recovery with Multi-Agent GAIL
To recover stationary policies in the DEC-POMDP, we use MA-GAIL with $n$ individual discriminators $\mathbf{D}=\{D_{1},D_{2},...,D_{n}\}$ and match occupancy measures on transformed expert demonstrations. For the generator network ${\pi}$, we use Multi-Agent Proximal Policy Optimization (MAPPO) with parameter sharing in which all learners use a single policy network. This applies to our environment since our agents have identical observation and action spaces (homogeneous agents) [29]. Using individual discriminators ensures that each learner strictly matches the occupancy measures of a particular expert. For homogeneous agents, however, this prevents generalization as learners receive poor feedback if they behave like any other experts. We address this through expert demonstration sharing and allow individual discriminators to compare trajectories from their learners with all expert demonstrations available. This ensures that learners are positively rewarded for demonstrating any valid expert behavior instead of the particular behavior from one expert.
@@ V EXPERIMENT
We evaluate the performance of our proposed approach in three classical swarm robotic scenarios: aggregation, homing, and obstacle avoidance. We model these scenarios as practical cooperative and single-objective in a shared environment. We consider a swarm size of $3$ and represent our swarming agents as unmanned aerial vehicles (UAVs) with inaccessible controllers. To improve learning representation, we reduce the complexity of the shared environment and separate it into motion and control layers. The control layer is a discretized grid world representation of the shared environment with reduced state and action space. We compare the performance of the proposed approach with behavior cloning (BC), where a direct mapping between expert states and actions is learned [30], and PS-AIRL [15].
@@ V-A Swarm Scenarios
The objective in the Homing scenario differs from aggregation in that the clustering zone—home position—is fixed and cannot be dynamically chosen by swarming UAVs. In this scenario, the UAVs must explore the environment and locate the home positions before the episode ends. Once a UAV finds a home position, it must remain there until all other UAVs have homed. We model $r_{n}$ as the maximum cohesion between UAV $n$ and any home position at a given time step.
In a new behavior that we refer to as obstacle avoidance, UAVs must navigate the shared environment without interacting with fixed-location inactive UAVs in the shared environment. This scenario differs from existing obstacle avoidance scenarios in that the UAVs can access the positions already occupied by the inactive UAVs in the control layers. However, they receive a large negative reward for doing this. The motivation for this behavior is that it is crucial to maintain a safe distance from unknown entities in a practical shared environment, even if they seem inactive. The UAVs also receive a small negative reward for insufficient exploration. We model $r_{n}$ as a large constant $-c$ when the cohesion between the UAV $n$ and any fixed position UAV is 00 and 00 otherwise.
@@ V-B Simulation Environment
We implement the motion layer for the shared environment in Webots [31] using simulated Crazyflies 2.0 drones [32] as the swarming UAVs. The simulation boundary is a $3$m by $3$m continuous rectangular world. UAVs can move in all directions and can detect obstacles using onboard range sensors. Figure 1 shows the position trace of the experts, learners, and random swarming UAVs in the obstacle avoidance scenario between $t=0$ and $t=300s$ with small red crosses. The three larger crosses on each setup mark the initial positions.
The control layer is a $10\times 10$ grid world environment. Here, the action space in the is limited to $a\in\mathbb{R}^{5}$ corresponding only to the high-level control of the UAVs {stop, right, left, forward, and backward}. Low-level motion controls such as lift-off, turning, translation, and hovering are implemented deterministically in the motion layer. Agents in the grid world can observe the positions of other entities up to $6$ grid positions in all directions. Given that only two fixed-position entities are in the aggregation scenario, the agent’s observation space is $o\in\mathbb{R}^{10}$ in this scenario and $o\in\mathbb{R}^{12}$ in others. All episodes run for a fixed duration of $50$ time steps in the control layer. This corresponds to an episode duration of about $300s$ in the motion layer.
@@ V-C1 Expert Demonstrations
To generate expert demonstrations of collective behavior in each scenario, we solve the corresponding Dec-POMDP grid world using PS-MAPPO for $100,000$ training episodes. After training, we generate an expert demonstration data pool of $1,000$ trajectories in each scenario. We also generate noisy expert demonstrations by varying expert optimality $\epsilon\in[0,1]$, where 00 implies optimal experts and $1$ implies experts sampling actions at random. It should be noted that trajectories in the expert demonstrations data pool are randomly generated and may contain similar expert UAV behavior.
V-C2 Learners UAVs
@@ VI RESULTS
We demonstrate the advantage of our behavior reconstruction algorithm in two different ways. In the first setup, we initialize learner UAVs from starting positions present in $\mathcal{D}$ for every evaluation episode. Figure 2 shows the performance comparison between the proposed approach (MA-GAIL), PS-AIRL, and BC trained using $400$ expert demonstrations over $200$ evaluation episodes with unnormalized reward values in all scenarios. As the figure demonstrates, our approach closely reproduces expert behavior in all scenarios compared to BC and PS-AIRL. We attribute this impressive performance across distinct swarm scenarios to the transformed expert demonstrations that sufficiently describe the multi-agent interactions in the shared environment and expert demonstration sharing which increases the set of valid expert behaviors. PS-AIRL outperformed BC in aggregation and homing scenarios but failed to maintain its superiority in the obstacle avoidance scenario. We attribute this to the sparsity of the reward function in the obstacle avoidance scenario, which forces the experts to conservatively explore a small area in the shared environment and avoid the large negative rewards. The abundance of sequential data from this region makes it easy for BC to clone expert actions when initialized from starting positions close to the area and outperform PS-AIRL. Conversely, the continuous reward function in aggregation and homing scenarios provides a normal distribution of state-action pairs in expert demonstrations making it difficult for BC to clone expert actions. It should be noted that evaluation results are from learners trained using $400$ expert demonstrations as they represent the best-performance region for PS-AIRL and BC.
@@
Transforming expert demonstration to generate informed and explainable multi-agent interactions improves learning representation and facilitates accurate behavior reconstruction. Furthermore, using $n$ individual discriminators while allowing them to share the transformed demonstrations reduces the complexity of each discriminator network and guarantees learners will be intuitively rewarded for all valid expert behavior. Nevertheless, these improvements introduce certain limitations. First, cohesion may not sufficiently describe multi-agent interactions in complex swarm scenarios involving multiple collective behaviors, and the search for a suitable interaction quantifier in such scenarios might not be trivial. A straightforward approach to address this is to transform expert demonstrations using several interaction metrics as in [16]. However, this may introduce redundancy and increase the computation budget as the swarm size increases. Second, using $n$ discriminators and sharing expert demonstrations between them can introduce scalability issues as $n$ becomes very large. While we address this through distributed discriminator training in this work, several discriminators (e.g. $n>100$) may be computationally expensive to train even in parallel. Finally, the challenge of collecting an adequate number of expert demonstrations remains, and we have not optimized our model to use minimal demonstrations. We intend to focus on these limitations in future works.
@@ VII CONCLUSION
In this work, we reconstructed collective behavior in shared environments without having access to the swarm controller. We achieve this by transforming expert demonstrations into state features that sufficiently describe the multi-agent interactions between entities in the shared environments. We investigate three distinct classical swarm robotics modeled as practical scenarios using UAVs. To improve learning representation, we separate the shared environments into motion and control layers and model the control layers as DEC-POMDPs grid world environments. Our results in the control layer demonstrate the capability of our MA-GAIL approach to accurately reconstruct observed collective behaviors in practical scenarios compared to existing reconstruction algorithms. We show that transforming expert demonstrations into state features that sufficiently describe multi-agent interactions and allowing learner robots to behave like any expert can facilitate swarm behavior prediction in distinct or unrelated swarm scenarios. In the future, we will investigate behavior reconstruction and recognition in complex practical scenarios involving multiple collective behaviors. As cohesion may not sufficiently describe the multi-agent interactions in these scenarios, we will investigate the automated discovery of explainable interaction quantifiers to achieve this. Future work will also consider the real-world constraints in experiments with physical multi-robot systems.

### v2

@@ Abstract
Understanding collective behavior and how it evolves is important to ensure that robot swarms can be trusted in a shared environment. One way to understand the behavior of the swarm is through collective behavior reconstruction using prior demonstrations. Existing approaches often require access to the swarm controller which may not be available. We reconstruct collective behaviors in distinct swarm scenarios involving shared environments without using swarm controller information. We achieve this by transforming prior demonstrations into features that describe multi-agent interactions before behavior reconstruction with multi-agent generative adversarial imitation learning (MA-GAIL). We show that our approach outperforms existing algorithms in spatial organization, and can be used to observe and reconstruct a swarm’s behavior for further analysis and testing, which might be impractical or undesirable on the original robot swarm.
@@ I INTRODUCTION
Swarm robotics and its applications are transitioning from the laboratory to the real world [1, 2], and it is expected to lead to the large-scale deployment of multiple robots in environments that are shared between robots and humans. For the robot-robot and human-robot interactions to be seamless, robot swarms must be safe and trustworthy [3]. Ensuring that a swarm is safe and trustworthy in shared environments requires precise and continuous knowledge of their collective behavior and how it evolves. In scenarios where swarm controllers are directly accessible, collective behavior can be modeled using the controllers [4]. However, in practical scenarios, swarm controllers may not be accessible for various reasons (e.g., inability to extract controllers from natural swarms). In robot swarms, access to controllers may be restricted or impractical due to various reasons such as encryption of controller information due to strategic or privacy concerns [5, 6]. To this end, understanding the collective behavior of a swarm and modeling its dynamics without using swarm controller information requires thorough research.
@@
Collective behavior reconstruction and recognition are two established methods for modeling swarm dynamics or explaining collective behavior [7]. Collective behavior reconstruction can be model-based or data-driven. In model-based approaches, behavior reconstruction is achieved using a mathematical or regression model [8, 9]. In data-driven approaches, multi-agent interactions are statistically extracted or learned from prior demonstrations to reproduce observed behavior. Recently, data-driven approaches used imitation learning (IL) algorithms such as inverse reinforcement learning (IRL) and generative adversarial imitation learning (GAIL) for improved reconstruction accuracy in swarm scenarios modeled as multi-agent systems [10, 11, 12, 13, 14, 15]. Genetic programming and graphical neural networks have also been used for data-driven behavior reconstruction with swarm controllers extracted from video demonstrations [16] and swarm behavior prediction [17]. For recognition of collective behaviour, binary classification of observed behavior as defined or undefined collective behavior is a common approach [18, 19, 20]. This approach does not scale, as each swarm scenario requires a unique classifier. Multinomial classification addresses this and has been achieved in closely related swarming scenarios using predefined and learned multi-agent interaction quantifiers [21, 22, 23].
@@
Existing data-driven behavior reconstruction approaches extract multi-agent interactions from expert demonstrations without capturing swarm-environment interactions. As a result, the recovered multi-agent interactions cannot accurately reconstruct or predict expert behavior. While this issue is addressed in [17], their approach relied on learned extraction of multi-agent interactions and it is difficult to explain how the robots interact with the environment.
@@
In this work, we investigate the reconstruction of collective behavior in three spatial organization tasks involving shared environments without using swarm controller information. We consider three common swarm robotic scenarios: aggregation, homing, and obstacle avoidance. We model these scenarios as single-objective swarm scenarios where swarming agents interact with each other and the environment while completing the swarm objective. We generate informed and explainable multi-agent interactions through feature transformation of expert demonstrations and, achieve near-optimal behavior reconstruction using multi-agent GAIL. We show that even when learner robots are initialized from unforeseen states, they perform similarly to the expert robots in all investigated scenarios.
@@ II RELATED WORKS
IRL has been extensively used in the literature for the reconstruction of collective behaviour as it recovers the underlying reward functions while reproducing expert behavior. Šošić et al. [11] reconstructed the behavior of a homogeneous swarm by assuming that all agents are interchangeable and share a central reward function, thereby reducing the problem to a single-agent IRL. Another study [12] reconstructed the collective behavior observed in a flock of pigeons by recovering individual reward functions for each pigeon. This individualistic approach exposed the multi-agent interactions in the flock and allowed the researchers to model the leader-follower hierarchy. In [24], a similar individualistic reward function approach was used to evolve the robot controller using IRL by manually specifying the desired goal location or the path. Other studies [13, 10] extended the IRL framework to multi-agent IRL to simultaneously recover the reward functions of several agents. Liu et al. [14], however, integrated GAIL with population-based training for collective behavior reconstruction. Besides IRL and GAIL, other machine learning methods have been used to reconstruct collective behavior. In [16], genetic programming was used to extract explainable controllers from video demonstrations of collective behavior with $8$ predefined swarm interaction metrics defining the fitness measure. Zhou et al. [17] used graphical neural networks to imitate the behavior of expert robots and predict trajectories. Their approach considers swarm-environment interactions and modeled robots and environmental entities as graph nodes, but does not provide explainable interactions or controllers. Most of these works demonstrate behavior reconstruction in a single scenario or multiple similar scenarios (e.g., swarming and schooling). However, Yu et al. [15] used Adversarial Imitation Learning with parameter sharing (PS-AIRL) for behavior reconstruction in distinct swarming scenarios. Their approach focused on homogeneous biological swarms and did not consider swarm-environment interactions. They also require access to the original swarm controllers which is rarely available in practical scenarios. In contrast, our approach considers swarm-environment interactions and reconstructs expert behavior without accessing robot controllers. We also generate informed multi-agent interactions that can be used to explain swarm behavior.
@@ IV-B Policy Recovery with Multi-Agent GAIL
To recover stationary policies in the DEC-POMDP, we use MA-GAIL with $n$ individual discriminators $\mathbf{D}=\{D_{1},D_{2},...,D_{n}\}$ and match occupancy measures on transformed expert demonstrations. For the generator network ${\pi}$, we use Multi-Agent Proximal Policy Optimization (MAPPO) with parameter sharing in which all learners use a single policy network. This applies to our environment since our agents are homogeneous and have identical observation and action spaces [29]. Using individual discriminators ensures that each learner strictly matches the occupancy measures of a particular expert. However, this prevents generalization for homogeneous agents as learners receive poor feedback if they behave like any other expert. We address this through expert demonstration sharing and allow individual discriminators to compare trajectories from their learners with all expert demonstrations available. This ensures that learners are positively rewarded for demonstrating any valid expert behavior instead of the particular behavior from one expert.
@@ V EXPERIMENT
We evaluate the performance of our proposed approach in three classical swarm robotic scenarios: aggregation, homing, and obstacle avoidance. We model these scenarios as cooperative and single-objective in a shared environment. We consider a swarm size of $3$ and represent our swarming agents as uncrewed aerial vehicles (UAVs) with inaccessible controllers. To improve learning representation, we reduce the complexity of the shared environment and separate it into motion and control layers. The control layer is a discretized grid world representation of the shared environment with reduced state and action space. We compare the performance of the proposed approach with PS-AIRL [15] and behavior cloning (BC), where a direct mapping between expert states and actions is learned [30].
@@ V-A Swarm Scenarios
The objective in the homing scenario differs from aggregation in that the clustering zone—home position—is fixed and cannot be dynamically chosen by the robots. In this scenario, the UAVs must explore the environment and locate the home positions before the episode ends. Once a UAV finds a home position, it must remain there until all other UAVs have homed. We model $r_{n}$ as the maximum cohesion between UAV $n$ and any home position at a given time step.
In a new behavior that we refer to as obstacle avoidance, UAVs must navigate the shared environment without interacting with fixed-location inactive UAVs in the shared environment. This scenario differs from existing obstacle avoidance scenarios in that the UAVs can access the positions already occupied by the inactive UAVs in the control layers. However, they receive a large negative reward for doing this. The motivation for this behavior is that it is crucial to maintain a safe distance from unknown entities in a practical shared environment, even if they seem inactive. The UAVs also receive a small negative reward for insufficient exploration. We model $r_{n}$ as a large constant $-c$ when the cohesion between the UAV $n$ and any fixed position UAV is maximized and 00 otherwise.
@@ V-B Simulation Environment
We implement the motion layer for the shared environment in Webots [31] using simulated Crazyflies 2.0 drones [32] as the swarm. The simulation boundary is a $3$m by $3$m continuous rectangular world. UAVs can move in all directions and can detect obstacles using onboard range sensors.
The control layer is a $10\times 10$ grid world environment. Here, the action space in the is limited to $a\in\mathbb{R}^{5}$ corresponding only to the high-level control of the UAVs {stop, right, left, forward, and backward}. Low-level motion controls such as lift-off, turning, translation, and hovering are implemented deterministically in the motion layer. Agents in the grid world can observe the positions of other entities up to $6$ grid positions in all directions. Given that only two fixed-position entities are in the aggregation scenario, the agent’s observation space is $o\in\mathbb{R}^{10}$ in this scenario and $o\in\mathbb{R}^{12}$ in others. All episodes run for a fixed duration of $50$ time steps in the control layer. This corresponds to an episode duration of about $300s$ in the motion layer. Figure 1 shows the position trace of the experts, learners, and random UAV behavior in the obstacle avoidance scenario between $t=0$ and $t=300s$ with small red crosses. The three larger crosses on each setup mark the initial positions.
@@ V-C1 Expert Demonstrations
To generate expert demonstrations of collective behavior in each scenario, we train expert UAVs in the corresponding Dec-POMDP grid world using PS-MAPPO for $100,000$ training episodes. After training, we generate an expert demonstration data pool of $1,000$ trajectories in each scenario. We also generate noisy expert demonstrations by varying expert optimality $\epsilon\in[0,1]$, where 00 implies optimal experts and $1$ implies experts sampling actions at random. It should be noted that trajectories in the expert demonstrations data pool are randomly generated and may contain similar expert UAV behavior.
V-C2 Learner UAVs
@@ VI RESULTS
We demonstrate the advantage of our behavior reconstruction algorithm in two different ways. In the first setup, we initialize learner UAVs from starting positions present in $\mathcal{D}$ for every evaluation episode. Figure 2 shows the performance comparison between the proposed approach (MA-GAIL), PS-AIRL, and BC trained using $400$ expert demonstrations over $200$ evaluation episodes with unnormalized reward values in all scenarios. As the figure demonstrates, our approach closely reproduces expert behavior in all scenarios compared to BC and PS-AIRL. This high performance across distinct swarm scenarios can be attributed to the transformed expert demonstrations, which sufficiently describe the multi-agent interactions in the shared environment. It can also be attributed to expert demonstration sharing, which increases the set of valid expert behaviors, thus allowing learners to easily reproduce expert behaviors regardless of the scenario. PS-AIRL outperformed BC in aggregation and homing scenarios but failed to maintain its superiority in the obstacle avoidance scenario. We attribute this to the sparsity of the reward function in the obstacle avoidance scenario, which forces the experts to conservatively explore a small area in the shared environment and avoid the large negative rewards. The abundance of sequential data from this region makes it easy for BC to clone expert actions when initialized from starting positions close to the area and outperform PS-AIRL. Conversely, the continuous reward function in aggregation and homing scenarios provides a normal distribution of state-action pairs in expert demonstrations making it difficult for BC to clone expert actions. It should be noted that evaluation results are from learners trained using $400$ expert demonstrations as they represent the best-performance region for PS-AIRL and BC.
@@
Transforming expert demonstration to generate informed and explainable multi-agent interactions improves learning representation and facilitates accurate behavior reconstruction. Furthermore, using $n$ individual discriminators while allowing them to share the transformed demonstrations reduces the complexity of each discriminator network and guarantees learners will be intuitively rewarded for all valid expert behavior. Nevertheless, these improvements introduce certain limitations. First, cohesion may not sufficiently describe multi-agent interactions in complex swarm scenarios involving multiple collective behaviors, and the search for a suitable interaction quantifier in such scenarios might not be trivial. A straightforward approach to address this is to transform expert demonstrations using several interaction metrics as in [16]. However, this may introduce redundancy and increase the computation budget as the swarm size increases. Second, using $n$ discriminators and sharing expert demonstrations between them can introduce scalability issues as $n$ becomes very large. While we address this through distributed discriminator training in this work, several discriminators (e.g., $n>100$) may be computationally expensive to train even in parallel. Finally, the challenge of collecting an adequate number of expert demonstrations remains, and we have not optimized our model to use minimal demonstrations. We intend to focus on these limitations in future works.
@@ VII CONCLUSION
In this work, we reconstructed collective behavior in shared environments without having access to the swarm controller. We achieve this by transforming expert demonstrations into state features that sufficiently describe the multi-agent interactions between entities in the shared environments. We investigate three distinct classical swarm robotics. To improve learning representation, we separate the shared environments into motion and control layers and model the control layers as DEC-POMDPs grid world environments. Our results in the control layer demonstrate the capability of our MA-GAIL approach to accurately reconstruct observed collective behaviors in spatial organisation (i.e., aggregation, goal homing, and obstacle avoidance) compared to existing reconstruction algorithms. We show that transforming expert demonstrations into shared state features that sufficiently describe multi-agent interactions improves behavior reconstruction accuracy in distinct or unrelated swarm scenarios. In the future, we will investigate behavior reconstruction and recognition in complex practical scenarios involving multiple collective behaviors. As cohesion does not sufficiently describe all multi-robot interactions in more complex scenarios, we will investigate the automated discovery of explainable interaction quantifiers to achieve this. Future work will also consider the real-world constraints in experiments with physical multi-robot systems.